OpenStudy (eskelle):

What values for theta(0

OpenStudy (eskelle):

this is the rest of the problem, just too much to type out

OpenStudy (hannahaz):

what values for theta(0 less than or equal to theta less than or equal to 2 pi) satisfy the equation? 3sin theta = sin theta -1 3sinx=sinx-1 3sinx-sinx=-1 2sinx=-1 sinx=-1/2 x=7π/6,11π/6 (in quadrants III and IV where sin<0)

OpenStudy (hannahaz):

2sinθcosθ + √2cosθ = 0 , 0 ≤ θ ≤ 2π cosθ(2sinθ + √2) = 0 So, cosθ = 0 or sinθ = -√2/2 If cosθ = 0, then θ = π/2, 3π/2 If sinθ = -√2/2, then θ = 5π/4, 7π/4

OpenStudy (hannahaz):

Let x = theta 2sin(x)cos(x) + cos(x)√2 = 0 Factor out cos(x). cos(x)(2sin(x) + √2) = 0 Set cos(x) equal to zero. cos(x) = 0 x = π/2 x = 3π/2 The solutions are x = π/2 and x = 3π/2

OpenStudy (hannahaz):

hope that helps

OpenStudy (eskelle):

those arent options

jimthompson5910 (jim_thompson5910):

@hannahaz has solved the first half you also need to solve \(\Large \sin(\theta) + \sqrt{2} = 0\) to get the other two solutions

OpenStudy (eskelle):

what is theta equal to again

jimthompson5910 (jim_thompson5910):

oh nvm, @hannahaz wrote this out `If sinθ = -√2/2, then θ = 5π/4, 7π/4`

OpenStudy (eskelle):

but i dont see that as an option? unless i am just blind what letter is it

OpenStudy (eskelle):

all the options have 4

jimthompson5910 (jim_thompson5910):

well you combine the 2 sets of values to get these four values total π/2, 3π/2, 5π/4, 7π/4

jimthompson5910 (jim_thompson5910):

each factor will have 2 solutions there are 2 factors so 2*2 = 4 solutions total

OpenStudy (eskelle):

ok so thats the final answer?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (eskelle):

thank you

jimthompson5910 (jim_thompson5910):

np